Find an equation of the line that has a slope of 10 and a y intercept of (0,8). Write your answer in the form y = Mx + b

Answer by ikleyn(53751)

.
Find an equation of the line that has a slope of 10 and a y intercept of (0,8).
Write your answer in the form y = Mx + b
~~~~~~~~~~~~~~~~

If the slope is m and y-intercept is the point (0,b),


then a straight line equation of the form y = Mx + b is


    y = mx + b,  or, in your case,  y = 10x + 8.    ANSWER

Solved, answered and explained.

-----------------

To see other similar solved problems (your future TEMPLATES), look into the lesson
    - Equation for a straight line having a given slope and passing through a given point
in this site.

Learn the subject from there.